OpenStudy (anonymous):

Find all polar coordinates of point P where P = ordered pair (5 , pi/3).

OpenStudy (anonymous):

hi :DD

OpenStudy (anonymous):

the radial lenghth is five and the angle from the x axis is pi/3..is'nt this already in polar co ordinates?

OpenStudy (anonymous):

so, obviously, the ordered pair (5, pi/3) works, since that is P. But if you go another 2pi around, you have the ordered pair (5, 7pi/3) which is the same point. in fact, if you find the ordered pair \[(5, (2\pi n)+\pi/3)\] for any integer n, it gets you to the same point

OpenStudy (anonymous):

but there's one more thing to consider...

OpenStudy (anonymous):

the point (-5, -2pi/3) also works...as seen in the diagram below 

OpenStudy (anonymous):

so, the points \[(-5, (2\pi n)-2\pi/3)\] also work.

OpenStudy (anonymous):

(5, pi divided by 3 + nπ) or (-5, pi divided by 3 + nπ) (5, pi divided by 3 + (2n + 1)π) or (-5, pi divided by 3 + 2nπ) (5, pi divided by 3 + 2nπ) or (-5, pi divided by 3 + (2n + 1)π) (5, pi divided by 3 + 2nπ) or (-5, pi divided by 3 + 2nπ)

OpenStudy (anonymous):

Those are my options

OpenStudy (anonymous):

so based on my first set, you can narrow it down to one of the last two, right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so of those 2, which seems more likely? if you don't know, it's ok, i just want to see if you can reason it out....

OpenStudy (anonymous):

Last one?

OpenStudy (anonymous):

so let's consider tha last one...specifically, the second set of the last one, (-5, pi/3 + 2npi) look at the circle i drew and imagine you are standing on the center... you turn so that you are facing pi/3. Now, you walk backwards 5 (walk along the line until you hit the circle...where would you be?

OpenStudy (anonymous):